Please help me solve this equation, step by step with websites suggestions were I can practice more. i would really appreciate it. Thank you.
If r = 5 z then 15 z = 3 y, then r =

Answer by Theo(13342)

i usually solve these by finding a common variable and making everythign in terms of that common variable.

your equations are:

r = 5z
15z = 3y

the common variable appears to be z.

solve for z in both equations to get:

z = r/5
z = 3y/15

since r/5 = z and 3y/15 = z, then r/5 = 3y/15.

solve for r as follows:

r/5 = 3y/15
multiply both sides of the equation by 5 to get:
r = 3y/15 * 5 = y/5 * 5 = y
the equation simplifies to:
r = y

you can also just solve for one of the variables in terms of the other and then substitute for that variable.

start with:
r = 5z
15z = 3y

solve for z in terms of y to get:
z = 3y/15 = y/5

replace z with y/5 in the first equation to get:
r = 5 * y/5 which becomes:
r = y